A ball is thrown horizontally at a velocity of 33.0 m/s from a cliff 24.0 meters high. How far from the base of the cliff does the ball land?



1
Expert's answer
2020-12-15T11:47:33-0500

Let's first find the time that the ball takes to reach the ground:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 24\ m}{9.8\ \dfrac{m}{s^2}}}=2.21\ s."

Finally, we can find the distance from the base of the cliff to the place where the ball lands:


"x=v_0t=33.0\ \dfrac{m}{s}\cdot 2.21\ s=73\ m."

Answer:

"x=73\ m."